# Hallucination to Truth: A Review of Fact-Checking and Factuality Evaluation in Large Language Models

### v1

@@ 4.4.3  Fine-tuning Architectures for Optimizing Fact-checking Performance
Fact-checking performance is primarily optimized through two primary approaches: (1) the development and fine-tuning of specific model architectures, and (2) the design of advanced prompting strategies for LLMs. These methods are often combined within complex fact-checking pipelines.
Fine-tuning smaller transformer models on synthetic data. This approach fine-tunes pre-trained transformer models on structured synthetic data, often combined with standard entailment datasets, to teach them nuanced fact-checking against grounding documents [10]. The strategy centers on generating challenging training instances that help models verify atomic facts across multiple sentences, with models like MiniCheck-FT5, RBTA, and DBTA outperforming larger LLMs such as GPT-4 in specific benchmarks like LLM-AGGREFACT [10]. Notably, MiniCheck-FT5 achieves a 4.3% improvement over AlignScore using a significantly smaller dataset. Difficulty aggregating evidence and reasoning over multiple facts, the method mitigates these through targeted synthetic data and simple aggregation strategies like majority voting [11].
Fine-tuning small language models (SLMs) for task-specific performance. This method focuses on fine-tuning SLMs for task-specific applications like fake news detection. The strategy includes training SLMs directly on the target dataset and leveraging LLM-generated rationales via architectures like the ARG network, with a distilled version (ARG-D) for efficiency [12]. Fine-tuned BERT models have outperformed GPT-3.5 in fake news detection, and ARG/ARG-D exceed baseline methods combining both SLM and LLM capabilities [12]. LLMs’ difficulty in fully utilizing their reasoning for domain-specific tasks and their inability to fully replace SLMs in these contexts. Mitigations involve using LLMs as rationale providers to guide SLMs and exploring advanced prompting and model combinations for better results [12].
Instruction Fine-tuned LLMs as Verifiers within a Pipeline Framework. This framework integrates an instruction fine-tuned LLM as a verifier within a fact-checking pipeline, where the LLM evaluates claims based on contextually retrieved evidence. The process includes claim atomization using Mistral-7B, relevance-based evidence retrieval and re-ranking, and final inference by the LLM, producing interpretable credibility reports [42]. The "Yours Truly" framework achieves a 94% F1-score, significantly outperforming other systems, with the Claim Atomizer alone boosting performance from 64% to 93% [42].
@@ 4.4.4  Domain-specific Training for Model Adaptation in Specialized Knowledge Areas
Domain-specific adaptation and fine-tuning are highlighted as crucial strategies for enhancing the performance and reliability of models, particularly LLMs, in automated fact-checking within specialized knowledge areas. The need for such domain specificity stems from the observation that LLM factual accuracy and vulnerability to hallucinations can vary significantly across different domains. While general-purpose models may perform well in broad areas, models fine-tuned or adapted for specific domains, such as medicine or science, often demonstrate superior performance in those particular fields. Sources discuss the application of fact-checking techniques across various specialized domains, including medical/biomedical, political, scientific, law, general biographic, and news [14, 30, 28, 31, 68].
Several approaches are explored for achieving domain adaptation in fact-checking systems. One involves fine-tuning smaller transformer models on target datasets specific to a domain or task within fact-checking, such as claim detection or veracity prediction, which has shown surprising efficacy and can outperform larger, general LLMs in specific contexts [48, 11]. Another approach directly involves fine-tuning LLMs themselves or using techniques like instruction-tuning on domain-relevant data or tasks [30, 48, 31, 46, 42, 40]. Frameworks like OpenFactCheck are proposed to allow users to customize fact-checkers for specific requirements, including domain specialization [68].
@@ 4.4.5  Comparative Summary and Trends
Prompt engineering has also progressed beyond simple Chain-of-Thought (CoT). The rise of organized hierarchical prompting techniques, such as HiSS, which break down complicated claims into verifiable steps to reduce hallucinations and improve reasoning, is a definite trend [21, 38]. Furthermore, domain-specific adaptation is becoming more and more important. Using specialized datasets and knowledge bases, models are being particularly trained or prompted for difficult domains such as news, climate science, and medicine to increase accuracy when general knowledge is inadequate [49, 76].
Integrating external domain-specific knowledge through methods like RAG or incorporating Knowledge Graphs is also a significant strategy to augment models with the necessary specialized information, sometimes in conjunction with fine-tuning or adaptation [36, 30, 15, 46, 18]. For example, the LEAF approach enhances medical question answering by integrating fact-checking results into RAG or using fact-checks for supervised fine-tuning [30]. Evaluation datasets tailored to specific domains or types of claims, such as SciFact for scientific claims, Climate-FEVER for climate science, EXPERTQA spanning multiple fields, or medical QA datasets, are utilized to benchmark the effectiveness of these domain-adapted systems [10, 14, 30, 28, 76].
@@ 4.5  Integration of RAG in Fact-Checking (RQ5)
RAG is a hybrid framework designed to enhance LLMs by integrating the retrieval of external knowledge into their generation process [56, 3]. This approach is pivotal for grounding LLM outputs in external evidence, thereby mitigating common issues such as hallucination, the generation of factually incorrect or nonsensical information, and reliance on potentially outdated internal knowledge [3, 15, 33]. Unlike methods solely based on fine-tuning the model’s internal parameters, RAG leverages external data sources, such as web search results or curated knowledge bases, to inform the LLM’s responses during inference [56, 15]. A significant advantage of RAG is its ability to produce responses that are not only more factually accurate and reliable but also offer improved transparency and credibility by providing explicit citations to the external sources used [48, 56]. This capability allows models to access and utilize current information, overcoming the limitations of knowledge cutoffs inherent in their training data [22]. An overview of the workflow of a basic RAG-based system is presented in Figure 10.
@@
In the domain of fact-checking, RAG systems play a crucial role in automating and improving the verification process [48]. LLM agents empowered with RAG can phrase search queries based on claims, retrieve relevant external contextual data, and utilize this information to assess the veracity of statements [48, 56]. This integration of external knowledge through RAG leads to enhanced accuracy in fact-checking, enabling the extraction of relevant evidence to support veracity predictions [48, 22, 53]. Approaches like Fact-Check-Then-RAG utilize the outcomes of fact-checking to refine the retrieval process itself and ensure that retrieved information specifically enhances factual accuracy [30]. RAG also supports more complex scenarios, including multimodal fact-checking, where it is used to extract both textual and image content and retrieve external information for reasoning [57, 22]. Furthermore, RAG-based systems can provide reasoned explanations for their verdicts, improving the interpretability of the fact-checking process [48, 14]. Variations like FFRR leverage fine-grained feedback from the LLM to optimize retrieval policy based on how well documents support factual claims [51]. While some approaches explore reducing reliance on external retrieval by leveraging the LLM’s internal knowledge, RAG is generally considered essential for effective fake news detection [57, 72].
@@
For inefficient and computationally expensive evidence retrieval, researchers have developed lightweight frameworks like Provenance [39], which use compact and open-source NLI models for verification instead of LLMs. FIRE [34], a framework that is designed for time and cost efficiency through an iterative process. Reinforcement retrieval models further show that a properly trained retriever does not add significant overhead during inference, as the costly feedback loop is only part of that training phase [51]. Due to the unavailability of evidence in specialized or low-resource domains, frameworks like LEAF [30], designed for the medical domain, use a specialized corpus such as MedRAG, which includes PubMed and textbooks, instead of relying solely on Google Search. The challenge of low-resource language fact checking is mitigated by translating the claims into high high-resource language, such as English, in some literature [48].
To avoid overwhelming the LLM context window and managing large volumes of retrieved framework like Provenance [39] use Relevancy Score combined with a TopK/TopP selection module to filter down to the most critical pieces of evidence before sending them to the verifier, LLM-AUGMENTER features a "Knowledge Consolidator" that prunes irrelevant information and synthesizes the remaining evidence into concise chains [51], Reinforcement Retrieval also suggests that using a smaller, optimal number of documents (e.g., the top 3-4) is often more effective than using a larger set that may introduce noise [51]. For mitigating noise and irrelevant information introduced by standard RAG methods, Reinforcement Retrieval addresses this directly using reinforcement learning to train the retriever; feedback from the LLM verifier acts as a reward signal, teaching the retriever to select more useful and factually relevant documents [51]. The LEAF framework uses a "Fact-Check-Then-RAG" approach, where an initial fact-check on the LLM’s output is used to guide a more targeted and accurate retrieval process [30].
Furthermore, for handling complex reasoning, conflicting evidence, and insufficient information, the PACAR framework employs a dynamic planner that can deploy tailored agents for tasks like numerical reasoning and entity disambiguation [14]. The LoCal framework is a multi-agent system specifically designed to handle the complexities of logical and causal fact-checking. For scenarios with conflicting or insufficient evidence [52], the AVERITEC dataset was created with a "Conflicting Evidence" label and a structured question-answer format that can represent evidential disagreements, providing a basis for training models to better navigate such ambiguity [52].
@@ 4.5.1  Comparative Summary and Trends
Figure 11: Limitations in RAG-based fact-checking and corresponding mitigation Strategies
@@ 5  Discussion
Our review explored the rapid adoption of LLMs in the complex task of automated fact checking. By examining a wide range of current studies, we have mapped out how we measure their success, the persistent problem of models generating false information (hallucinations), and the essential role and limitations of the datasets they rely on. We also looked into methods for improving LLMs, from prompt engineering and fine-tuning to the increasingly vital use of RAG. The research landscape reveals a field that is buzzing with innovation and showing great promise. It simultaneously highlights significant and complex hurdles that remain. If LLMs are to become truly reliable tools in the global effort against misinformation, these challenges demand ongoing and rigorous investigation [3, 48].
@@
Evaluation metrics. In the evaluation metrics domain (RQ1) for LLM-based fact checking, we can see the clear transition in classification scores towards more sophisticated, holistic and context-aware frameworks. The rise of rigorous benchmarks such as LLM-AGGREFACT [104] and the AVERITEC dataset [52], along with new centralized tools such as OpenFactCheck [44], marks a major step towards creating shared ways to measure how accurate LLMs are. However, even with this progress, many important issues remain unresolved. Furthermore, making models tough enough to handle misinformation that adapts, often as an adversary, including claims with subtle edits or those that change over time, remains a major hurdle [40]. The lack of strong methods for clear explanation and adaptation to new types of risks makes it harder for everyday users to judge whether the answers from LLMs are true. This shortfall also weakens the usefulness of these models in fast-changing real-life situations.
@@
Two key challenges are now drawing more attention: teaching models how to check their work and catch their own mistakes [29], and finding reliable ways to measure how closely their responses match the source material [50]. Despite advances in automated and LLM-driven assessments, Human Evaluation remains essential for nuanced aspects like explanation quality and contextual appropriateness [73], although it is resource intensive. In general, while progress is evident, significant lacunae persist. There is a pressing need for standardized metrics that robustly assess the quality of LLM-generated explanations [31], the resilience of the model against evolving misinformation [40], and the logical integrity of reasoning pathways [17]. These developments underscore an urgent and ongoing need for metrics that can holistically evaluate not only the veracity of claims, but also the provenance of supporting evidence [39] and the logical integrity of the LLM’s reasoning process.
@@
Hallucination in LLMs. The tendency of LLMs to hallucinate (RQ2), that is, to generate outputs that are linguistically fluent and coherent yet factually misleading or entirely unsubstantiated, remains a significant barrier to their trustworthy deployment in sensitive, high-stakes applications such as fact-checking [3, 21]. RAG has become a foundational mitigation strategy, designed to ground LLM responses in verifiable external knowledge and reduce the models’ reliance on their internal, and potentially flawed or outdated, parametric knowledge [48, 15, 107, 46]. New efforts to boost how truthful language models are include adding systems that automatically give feedback, helping refine answers over multiple tries [15]. Tools like Self-Checker are also being built—these are smart correction modules that let the model review and fix its own mistakes [20]. On top of that, researchers are testing ways to bring together evidence from several sources, with models like MEDICO showing how this kind of fusion can work [49]. This persistence of hallucinations is partly due to the inherent "black box" nature of current LLMs, the difficulty in comprehensively modeling nuanced world knowledge, and the escalating sophistication of adversarial attacks designed to exploit model vulnerabilities [50].
@@
Datasets for Fact-Checking. The datasets (RQ3) utilized for the training, fine-tuning, and rigorous evaluation of fact-checking LLMs are pivotal to their ultimate performance and generalizability. While foundational datasets such as FEVER [86] have been instrumental in catalyzing early research, the field is increasingly recognizing the necessity for more specialized, challenging, and contextually rich benchmarks. Illustrative examples include CliniFact, which is tailored for claims within the domain of clinical research [28], AVERITEC, with its distinct emphasis on real-world claims that necessitate web-based evidence retrieval [52], and BINGCHECK, specifically designed for assessing the factuality of LLM-generated text [20]. The intrinsic characteristics of these datasets—including their composition, scale, annotation quality, topical diversity, and potential inherent biases (e.g., political, cultural, or temporal) profoundly influence model performance, the ability of models to generalize to unseen domains or claim structures, and, consequently, the perceived effectiveness of various fact-checking methodologies [85, 54]. The development and meticulous curation of robust multilingual datasets [32] also represent a critical frontier for advancing the global applicability and equity of LLM-based fact-checking technologies.
@@
Optimization Strategies and Domain-specific Training. Prompt engineering, fine-tuning strategies, and domain-specific training (RQ4) have been shown to significantly modulate the efficacy of LLMs in complex fact-checking tasks. Advanced prompting techniques, such as the hierarchical step-by-step verification methods [21] or structured reasoning frameworks like PACAR [14] and BiDeV [37], frequently demonstrate superior outcomes when compared to simpler, more direct prompting approaches. The application of zero-shot and few-shot learning paradigms is also being actively explored for a range of related tasks, including claim matching [41, 47, 80], indicating a strong potential for efficient adaptation of LLMs with limited task-specific data. Furthermore, the surprising efficacy of smaller, specifically fine-tuned transformer models in certain fact-checking contexts [11] compellingly suggests that model scale is not the sole, nor always the primary, determinant of performance. This finding challenges the prevailing "bigger is better" narrative in LLM development and highlights the potential for more resource-efficient, specialized models to achieve competitive or even superior performance in targeted fact-checking applications, particularly when data and computational budgets are constrained. Domain-specific adaptations, which may involve the utilization of LLM-predicted credibility signals [55] or the development of highly specialized systems for critical domains such as medicine [38], are proving essential for achieving the nuanced understanding and reliable fact verification required in these contexts.
@@
The Integration of RAG. Incorporating LLMs with RAG (RQ5) is increasingly recognized as a central and indispensable strategy for enhancing the factuality of LLM outputs. This is achieved by providing models with dynamic access to external, often real-time, knowledge sources, thereby augmenting their inherent capabilities [48, 15, 107, 46]. Frameworks such as FIRE [34] are being developed to optimize iterative retrieval and verification processes, aiming for greater efficiency and accuracy. Other lines of research explore the application of reinforcement learning techniques to refine and optimize retrieval strategies [21]. Nevertheless, significant challenges persist within the RAG pipeline. These include the efficient and precise retrieval of truly relevant evidence from vast, heterogeneous, and often noisy information spaces; the effective fusion of information derived from multiple, potentially contradictory, sources [14]. Extending RAG to effectively handle multi-modal inputs [22, 57] and optimizing the timing and contextual relevance of information recommended by RAG-empowered agents [61] remain active and critical areas of ongoing investigation.

### v2

@@ 4.4.3  Fine-tuning Architectures for Optimizing Fact-checking Performance
Fact-checking performance is primarily optimized through two primary approaches: (1) the development and fine-tuning of specific model architectures and (2) the design of advanced prompting strategies for LLMs. These methods are often combined within complex fact-checking pipelines.
Fine-tuning smaller transformer models on synthetic data. This approach fine-tunes pre-trained transformer models on structured synthetic data, often combined with standard entailment datasets, to teach them nuanced fact-checking against grounding documents [10]. The strategy focuses on generating challenging training instances that help models verify atomic facts across multiple sentences, with models such as MiniCheck-FT5, RBTA, and DBTA outperforming larger LLMs such as GPT-4 in specific benchmarks like LLM-AGGREFACT [10]. Notably, MiniCheck-FT5 achieves a 4.3% improvement over AlignScore using a significantly smaller dataset. Difficulty aggregating evidence and reasoning over multiple facts, the method mitigates these through targeted synthetic data and simple aggregation strategies like majority voting [11].
Fine-tuning SLMs for task-specific performance. This method focuses on fine-tuning SLMs for task-specific applications like fake news detection. The strategy includes training SLM directly on the target dataset and using LLM-generated rationales via architectures like the ARG network, with a distilled version (ARG-D) for efficiency [12]. Fine-tuned BERT models have outperformed GPT-3.5 in fake news detection, and ARG/ARG-D exceed baseline methods that combine both SLM and LLM capabilities [12]. LLMs’ difficulty in fully utilizing their reasoning for domain-specific tasks and their inability to fully replace SLMs in these contexts. Mitigation strategies include employing LLMs as rationale providers to guide SLMs, as well as exploring advanced prompting techniques and model combinations to achieve improved performance [12].
Instruction Fine-tuned LLMs as Verifiers within a Pipeline Framework. This framework integrates an instruction-fine-tuned LLM as a verifier within a fact-checking pipeline, where the LLM evaluates claims based on contextually retrieved evidence. The process includes claim atomization using Mistral-7B, relevance-based evidence retrieval and re-ranking, and final inference by the LLM, producing interpretable credibility reports [42]. The "Yours Truly" framework achieves a 94% F1-score, significantly outperforming other systems, with the Claim Atomizer alone boosting performance from 64% to 93% [42].
@@ 4.4.4  Domain-specific Training for Model Adaptation in Specialized Knowledge Areas
Domain-specific adaptation and fine-tuning are highlighted as crucial strategies for enhancing the performance and reliability of models, particularly LLMs, in automated fact-checking within specialized knowledge areas. The need for such domain specificity arises from the observation that the factual accuracy and vulnerability of LLM to hallucinations can vary significantly between different domains. While general-purpose models may perform well in broad areas, models fine-tuned or adapted for specific domains, such as medicine or science, often demonstrate superior performance in those particular fields. Sources discuss the application of fact-checking techniques across various specialized domains, including medical/biomedical, political, scientific, law, general biographic, and news [14, 28, 30, 31, 44].
Several approaches are explored to achieve domain adaptation in fact-checking systems. One involves fine-tuning smaller transformer models on target datasets specific to a domain or task within fact-checking, such as claim detection or veracity prediction, which has shown surprising efficacy and can outperform larger, general LLMs in specific contexts [11, 48]. Another approach directly involves fine-tuning LLMs themselves or using techniques like instruction-tuning on domain-relevant data or tasks [30, 31, 40, 42, 46, 48]. Frameworks like OpenFactCheck are proposed to allow users to customize fact-checkers for specific requirements, including domain specialization [44].
@@ 4.4.5  Comparative Summary and Trends
Prompt engineering has also progressed beyond simple CoT. The rise of organized hierarchical prompting techniques, such as HiSS, which break down complicated claims into verifiable steps to reduce hallucinations and improve reasoning, is a definite trend [21, 38]. Furthermore, domain-specific adaptation is becoming more and more important. Using specialized datasets and knowledge bases, models are being particularly trained or prompted for difficult domains such as news, climate science, and medicine to increase the accuracy when general knowledge is inadequate [49, 75].
Integrating external domain-specific knowledge through methods like RAG or incorporating Knowledge Graphs is also a significant strategy to augment models with the necessary specialized information, sometimes in conjunction with fine-tuning or adaptation [15, 18, 36, 30, 46]. For example, the LEAF approach enhances medical question answering by integrating fact-checking results into RAG and using fact-checks for supervised fine-tuning [30]. Evaluation datasets tailored to specific domains or types of claims, such as SciFact for scientific claims, Climate-FEVER for climate science, EXPERTQA spanning multiple fields, or medical QA datasets, are utilized to benchmark the effectiveness of these domain-adapted systems [10, 14, 28, 30, 75].
@@ 4.5  Integration of RAG in Fact-Checking (RQ5)
RAG is a hybrid framework designed to enhance LLMs by integrating the retrieval of external knowledge into their generation process [3, 56]. This approach is pivotal for grounding LLM outputs in external evidence, thereby mitigating common issues such as hallucination, the generation of factually incorrect or nonsensical information, and reliance on potentially outdated internal knowledge [3, 15, 33]. Unlike methods solely based on fine-tuning the model’s internal parameters, RAG employs external data sources, such as web search results or curated knowledge bases, to inform the LLM’s responses during inference [15, 56]. A significant advantage of RAG is its ability to produce responses that are factually accurate and reliable, and offer improved transparency and credibility by providing explicit citations to the external sources used [48, 56, 117]. This capability allows models to access and utilize current information, overcoming the limitations of knowledge cutoffs inherent in their training data [22]. An overview of the workflow of a basic RAG-based system is presented in Figure 11.
@@
In the domain of fact-checking, RAG systems play a crucial role in automating and improving the verification process [48]. LLM agents empowered with RAG can phrase search queries based on claims, retrieve relevant external contextual data, and utilize this information to assess the veracity of statements [48, 56]. This integration of external knowledge through RAG leads to enhanced accuracy in fact-checking, enabling the extraction of relevant evidence to support veracity predictions [22, 48, 53]. Approaches like Fact-Check-Then-RAG utilize the outcomes of fact-checking to refine the retrieval process itself and ensure that retrieved information specifically enhances factual accuracy [30]. RAG also supports more complex scenarios, including multimodal fact-checking, where it is used to extract both textual and image content and retrieve external information for reasoning [22, 57]. Furthermore, RAG-based systems can provide reasoned explanations for their verdicts, improving the interpretability of the fact-checking process [14, 48]. Variations like FFRR utilize fine-grained feedback from the LLM to optimize retrieval policy based on how well documents support factual claims [51]. While some approaches explore reducing reliance on external retrieval by comprising the LLM’s internal knowledge, RAG is generally considered essential for effective fake news detection [57, 71].
Figure 12: Limitations in RAG-based fact-checking and mitigation Strategies. The red boxes illustrate the limitations, while the green boxes represent the corresponding mitigation strategies.
@@
For inefficient and computationally expensive evidence retrieval, researchers have developed lightweight frameworks like Provenance [39], which use compact and open-source NLI models for verification instead of LLMs. FIRE [34], a framework that is designed for time and cost efficiency through an iterative process. Reinforcement retrieval models further show that a properly trained retriever does not add significant overhead during inference, as the costly feedback loop is only part of that training phase [51]. Due to the unavailability of evidence in specialized or low-resource domains, frameworks like LEAF [30], designed for the medical domain, use a specialized corpus such as MedRAG, which includes PubMed and textbooks, instead of relying solely on Google Search. The challenge of fact-checking in low-resource languages is mitigated by translating the claims into high-resource languages, such as English, in some literature [48].
To prevent overwhelming the LLM’s context window and to efficiently manage large volumes of retrieved information, frameworks such as Provenance [39] employ a Relevancy Score in combination with TopK/TopP selection modules to filter the most critical evidence before it reaches the verifier. Similarly, LLM-AUGMENTER incorporates a "Knowledge Consolidator" that prunes irrelevant data and synthesizes the remaining evidence into concise reasoning chains [51]. Reinforcement Retrieval further demonstrates that limiting the number of documents, typically to the top three or four, can be more effective than including a larger set, which may introduce noise [51]. For mitigating noise and irrelevant information introduced by standard RAG methods, Reinforcement Retrieval addresses this directly using reinforcement learning to train the retriever; feedback from the LLM verifier acts as a reward signal, teaching the retriever to select more useful and factually relevant documents [51]. The LEAF framework uses a "Fact-Check-Then-RAG" approach, where an initial fact-check on the LLM’s output is used to guide a more targeted and accurate retrieval process [30].
Furthermore, to handle complex reasoning, conflicting evidence, and insufficient information, the PACAR framework employs a dynamic planner that can deploy tailored agents for tasks like numerical reasoning and entity disambiguation [14]. The LoCal framework is a multi-agent system specifically designed to handle the complexities of logical and causal fact-checking. For scenarios with conflicting or insufficient evidence [52], the AVERITEC dataset was created with a "Conflicting Evidence" label and a structured question-answer format that can represent evidential disagreements, providing a basis for training models to better navigate such ambiguity [52].
@@ 4.5.1  Comparative Summary and Trends
Table 7: Quantitative Synthesis of Key Findings on Fact-Checking Methodologies.
Claim & Core Insight Quantitative Evidence & Context Papers
Domain-Tuned SLMs Can Outperform Larger LLMs A fine-tuned BERT model achieved a +9.0% relative improvement in F1-score over the best-performing GPT-3.5-turbo configuration (76.5% vs. 70.2%) on the GossipCop dataset [12]. 4
A supervised ROBERTa-Base classifier outperformed zero-shot GPT-3.5-turbo by +9.6 absolute points in F1-macro on the FA-KES dataset (52.9% vs. 43.3%) [55].
A fine-tuned BioBERT for clinical claims achieved 80.2% accuracy on CliniFact, significantly outperforming both zero-shot (34.3%) and fine-tuned (53.6%) Llama3-70B [28].
MiniCheck-FT5 model (770M parameters), fine-tuned on specially generated synthetic data, achieved an average Balanced Accuracy of 74.7% on the LLM-AGGREFACT benchmark, which was statistically comparable to GPT-4’s score of 75.3 GPT-4’s score of 75.3% [10].
Advanced Prompting is Better than Standard CoT The HiSS (Hierarchical Step-by-Step) method surpassed vanilla CoT by +9.5 absolute points in F1-score on the RAWFC dataset (53.9% vs. 44.4%). 2
On the LIAR dataset, HiSS outperformed vanilla CoT by +7.1 points in F1-score (31.3% vs. 24.2%). HiSS also outperformed the more advanced ReAct agent framework by +4.1 absolute points on RAWFC [21].
At a detailed 5-class classification level (Level 2), the CLIMINATOR framework achieved an accuracy of 72.7%. This was substantially better than the GPT-4o advocate alone, which only achieved 56.5% accuracy on the same task[75].
RAG Provides Significant Performance Gains Providing external context (RAG) to GPT-4 on the PolitiFact dataset increased its accuracy on non-ambiguous verdicts from 75% to 89%. The accuracy on "true" claims jumped by +13.62 absolute points [48]. 5
The Fact-Check-Then-RAG method improved Llama 3 70B’s accuracy on the PubMedQA dataset from 60.60% to 73.60% (+13.0 absolute points) by using fact-checking results to guide retrieval [30].
An RAG pipeline using Mixtral achieved a 0.780 F1-score on the ’Refuted’ class on the Averitec development set [53].
On the RAWFC dataset, the fine-tuned FactLLaMA model with external knowledge achieved a Macro-F1 score of 0.5565. This significantly outperformed the same model fine-tuned without external knowledge, which only scored 0.5376 [46]
On the LIAR-RAW dataset, "Direct Prompting" of the LLM gave an F1 score of 27.0%. Simply adding a frozen (non-optimized) retriever in the FFRR-frozen setup increased the F1 score to 30.1%, demonstrating the inherent benefit of RAG. [51]
Hybrid and Multi-Agent Systems are More Effective Compared to strong LLM baselines like Flan-T5 and ChatGPT, the LoCal multi-agent system showed an average performance improvement of up to 7.75% in the gold evidence setting and up to 6.17% in the open book setting [16]. 4
The hybrid SLM+LLM ARG network improved F1-score over its BERT-only baseline by +3.1 absolute points on the Weibo21 dataset (78.4% vs. 75.3%) [12].
The PACAR framework, with specialized agents, outperformed a general ChatGPT baseline by +16.9 absolute points on HOVER 4-hop claims (72.61% vs. 55.72%) [14].
The FACT-AUDIT adaptive multi-agent framework demonstrated superior evaluation robustness over static, single-agent pipelines by dynamically assigning roles [45].
Automated Feedback Mechanisms Reduce Hallucinations The LLM-AUGMENTER system, using a BM25 knowledge consolidator and automated feedback, improved the KF1 score to 37.41 over the GPT KF1 score of 31.33 [15]. 4
The Self-Checker framework improved label accuracy on the BINGCHECK dataset from 21.0% (ReAct baseline) to 63.4% [20].
Medico’s multi-source evidence fusion and correction loop improved hallucination detection F1-score by +34.4 points over its baseline on the HaluEval dataset [49].
The Visual Fact Checker uses object detection and VQA models as automated "tools" to verify and correct initial caption proposals, significantly reducing hallucinations in detailed image captions [43].
Fine-tuning on Synthetic Data Boosts Performance The MiniCheck-FT5 model, trained on generated synthetic data, achieved a Balanced Accuracy of 74.7% on the LLM-AGGREFACT benchmark, a +4.3 absolute point improvement over the previous state-of-the-art. An ablation study showed that removing this synthetic data caused the model’s performance to drop by -14.8 absolute points [10]. 2
FACT-GPT was trained on a synthetic dataset of contradicting, entailing, or neutral claims generated by GPT-4, which enabled a smaller, specialized LLM to match the claim-matching accuracy of larger models [88].
@@ 5  Discussion
Our review explored the rapid adoption of LLMs in the complex task of automated fact-checking. By examining a wide range of current studies, we have mapped out how we measure their success, the persistent problem of models generating false information (hallucinations), and the essential role and limitations of the datasets they rely on. We also looked into methods for improving LLMs, from prompt engineering and fine-tuning to the increasingly vital use of RAG. The research landscape reveals a field that is buzzing with innovation and showing great promise. It simultaneously highlights significant and complex hurdles that remain. If LLMs are to become truly reliable tools in the global effort against misinformation, these challenges demand ongoing and rigorous investigation [3, 48].
@@
Evaluation metrics. In the evaluation metrics domain (RQ1) for LLM-based fact-checking, we can see the clear transition in classification scores towards more sophisticated, holistic, and context-aware frameworks. The rise of rigorous benchmarks such as LLM-AGGREFACT [10] and the AVERITEC dataset [52], along with new centralized tools such as OpenFactCheck [44], marks a major step towards creating shared ways to measure how accurate LLMs are. However, even with this progress, many important issues remain unresolved. Furthermore, making models tough enough to handle misinformation that adapts, often as an adversary, including claims with subtle edits or those that change over time, remains a major hurdle [40]. The lack of strong methods for clear explanation and adaptation to new types of risks makes it harder for everyday users to judge whether the answers from LLMs are true. This shortfall also weakens the usefulness of these models in fast-changing real-life situations.
@@
Two key challenges are now drawing more attention: teaching models how to check their work and catch their own mistakes [29], and finding reliable ways to measure how closely their responses match the source material [50]. Despite advances in automated and LLM-driven assessments, Human Evaluation remains essential for nuanced aspects like explanation quality and contextual appropriateness [72], although it is resource-intensive. In general, while progress is evident, significant lacunae persist. There is a pressing need for standardized metrics that robustly assess the quality of LLM-generated explanations [31], the resilience of the model against evolving misinformation [40], and the logical integrity of reasoning pathways [17]. These developments underscore an urgent and ongoing need for metrics that can holistically evaluate not only the veracity of claims, but also the provenance of supporting evidence [39] and the logical integrity of the LLM’s reasoning process.
@@
Hallucination in LLMs. The tendency of LLMs to hallucinate (RQ2), that is, to generate outputs that are linguistically fluent and coherent yet factually misleading or entirely unsubstantiated, remains a significant barrier to their trustworthy deployment in sensitive, high-stakes applications such as fact-checking [3, 21]. RAG has become a foundational mitigation strategy, designed to ground LLM responses in verifiable external knowledge and reduce the models’ reliance on their internal, and potentially flawed or outdated, parametric knowledge [15, 46, 48, 118]. New efforts to boost how truthful language models are include adding systems that automatically give feedback, helping refine answers over multiple tries [15]. Tools like Self-Checker are also being built. These are smart correction modules that let the model review and fix its own mistakes [20]. On top of that, researchers are testing ways to bring together evidence from several sources, with models like MEDICO showing how this kind of fusion can work [49]. This persistence of hallucinations is partly due to the inherent "black box" nature of current LLMs, the difficulty in comprehensively modeling nuanced world knowledge, and the escalating sophistication of adversarial attacks designed to exploit model vulnerabilities [50].
@@
Datasets for Fact-Checking. The datasets (RQ3) utilized for the training, fine-tuning, and rigorous evaluation of fact-checking LLMs are pivotal to their ultimate performance and generalizability. While foundational datasets such as FEVER [96] have been instrumental in catalyzing early research, the field is increasingly recognizing the necessity for more specialized, challenging, and contextually rich benchmarks. Illustrative examples include CliniFact, which is tailored for claims within the domain of clinical research [28], AVERITEC, with its distinct emphasis on real-world claims that necessitate web-based evidence retrieval [52], and BINGCHECK, specifically designed for assessing the factuality of LLM-generated text [20]. The intrinsic characteristics of these datasets, including their composition, scale, annotation quality, topical diversity, and potential inherent biases (e.g., political, cultural, or temporal) profoundly influence model performance, the ability of models to generalize to unseen domains or claim structures, and, consequently, the perceived effectiveness of various fact-checking methodologies [54, 95]. The development and meticulous curation of robust multilingual datasets [32] also represent a critical frontier to advance the global applicability and equity of LLM-based fact-checking technologies.
@@
Optimization Strategies and Domain-specific Training. Prompt engineering, fine-tuning strategies, and domain-specific training (RQ4) have been shown to significantly modulate the efficacy of LLMs in complex fact-checking tasks. Advanced prompting techniques, such as hierarchical step-by-step verification methods [21] or structured reasoning frameworks like PACAR [14] and BiDeV [37], frequently demonstrate superior outcomes when compared to simpler, more direct prompting approaches. The application of zero-shot and few-shot learning paradigms is also being actively explored for a range of related tasks, including claim matching [41, 47, 88], indicating a strong potential for efficient adaptation of LLMs with limited task-specific data. Furthermore, the surprising efficacy of smaller, specifically fine-tuned transformer models in certain fact-checking contexts [11] compellingly suggests that model scale is not the sole, nor always the primary, determinant of performance. This finding challenges the prevailing "bigger is better" narrative in LLM development and highlights the potential for more resource-efficient, specialized models to achieve competitive or even superior performance in targeted fact-checking applications, particularly when data and computational budgets are constrained. Domain-specific adaptations, which may involve the utilization of LLM-predicted credibility signals [55] or the development of highly specialized systems for critical domains such as medicine [38], are proving essential for achieving the nuanced understanding and reliable fact verification required in these contexts.
@@
The Integration of RAG. Incorporating LLM with RAG (RQ5) is increasingly recognized as a central and indispensable strategy for enhancing the factuality of LLM outputs. This is achieved by providing models with dynamic access to external, often real-time, knowledge sources, thereby augmenting their inherent capabilities [15, 46, 48, 118]. A framework such as FIRE [34] is being developed to optimize iterative retrieval and verification processes, aiming for greater efficiency and accuracy. Other lines of research explore the application of reinforcement learning techniques to refine and optimize retrieval strategies [21]. Nevertheless, significant challenges remain within the RAG pipeline. These include the efficient and precise retrieval of truly relevant evidence from vast, heterogeneous, and often noisy information spaces; the effective fusion of information derived from multiple, potentially contradictory, sources [14]. Extending RAG to effectively handle multimodal inputs [22, 57] and optimizing the timing and contextual relevance of information recommended by RAG-empowered agents [61] remain active and critical areas of ongoing investigation.